James pulls a sled with a mass of 80kg across his lawn. He pulls with 150N of force. a) Determine the force of friction if the acceleration is 0.5m/s2. b) What is the coefficient of friction?

 


1
Expert's answer
2021-11-04T10:16:40-0400

a)

"F-F_{fr}=ma,"

"F_{fr}=F-ma=150-80\cdot 0.5=110~N,"

b)

"F_{fr}=\mu N=\mu mg,"

"\mu=\frac{F_{fr}}{mg}=\frac{110}{80\cdot 9.8}=0.14."